# Predicting Case Suffixes With Activity Start and End Times: A Sweep-Line Based Approach

### v1

@@ Abstract
Predictive process monitoring techniques support the operational decision-making by predicting future states of ongoing cases of a business process. A subset of these techniques predict the remaining sequence of activities of an ongoing case (case suffix prediction). Existing approaches for case suffix prediction generate sequences of activities with a single timestamp (e.g. the end timestamp). This output is insufficient for resource capacity planning, where we need to reason about the periods of time when resources will be busy performing work. This paper introduces a technique for predicting case suffixes consisting of activities with start and end timestamps. In other words, the proposed technique predicts both the waiting time and the processing time of each activity. Since the waiting time of an activity in a case depends on how busy resources are in other cases, the technique adopts a sweep-line approach, wherein the suffixes of all ongoing cases in the process are predicted in lockstep, rather than predictions being made for each case in isolation. An evaluation on real-life and synthetic datasets compares the accuracy of different instantiations of this approach, demonstrating the advantages of a multi-model approach to case suffix prediction.
@@ 1 Introduction
Predictive Process Monitoring (PPM) techniques provide runtime insights to operational managers by exploiting models trained on event logs to predict the future states of ongoing cases of a process. For example, PPM techniques can predict the outcome of ongoing cases (e.g. will a customer accept or reject a product?) [21], the remaining time of ongoing cases [22], the next activity in a case, or the sequence of remaining activities in a case (a.k.a. the case suffix[16].
Existing approaches for case suffix prediction generate sequences of activities with a single timestamp (e.g., end timestamp). This is insufficient for use cases such as capacity planning and scheduling, where managers need both start and end timestamps to calculate processing and waiting times, assess resource utilization, and determine whether capacity is adequate for the expected workload.
@@
In this setting, we propose an approach for predicting case suffixes composed of activities with start and end timestamps. Our method employs a sweep-line-based technique [14], where suffixes for all ongoing cases are predicted collectively, rather than individually. The approach follows a three-stage prediction process. First, a model predicts the next activity. Given this activity, a second model estimates the inter-start time, i.e., the time elapsed between the start of the previous activity and the start of the predicted activity. Finally, a third model predicts the processing time i.e., the duration required to complete the activity once it has started. Using these predictions, we derive the start and end timestamps of each predicted activity instance in the case suffix. The start timestamp is obtained by adding the inter-start time to the start timestamp of the previous activity instance, while the end timestamp is derived by adding the processing time to the start timestamp of the predicted activity instance.
We report on an evaluation that compares the accuracy of our multi-model approach, along control-flow and temporal metrics, relative to approaches that predict the next activity and timestamp via a single model approach.
@@ 2 Related Work
Several studies have addressed the problem of predicting case suffixes via deep learning. Tax et al. [19] and Evermann et al. [6] proposed LSTM-based architectures to predict the next activity, the case suffix, and the remaining time until case completion, using one-hot and embedded sequence encodings, respectively. Camargo et al. [4] extended these approaches by proposing multi-output LSTM models capable of jointly predicting activity, end timestamp, and roles. Taymouri et al. [20] proposed a GAN-based approach, while Wuyts et al. [24] and Rama-Maneiro et al. [17] introduced transformer and attention-based encoder-decoder models to predict full suffixes in a single step. Ketykó et al. [10] highlight that suffix prediction performance is influenced by both architecture and sampling/decoding strategies. Rama-Maneiro et al. [15] propose DOGE, a DRL-based sampler that replaces traditional methods like Argmax with learned policies. Pasquadibisceglie et al. [12] proposed CNN-based case suffix predictors, and later extended this approach using Large Language Models (LLMs) [13].
@@
The above approaches follow a single model (SM) paradigm, which jointly predicts the next activity and one timestamp (either the start or the end time). None of these approaches is designed to estimate both the start and the end timestamps of the activities in the case suffix. In other words, none of these approaches sepatrates the waiting time and the processing time of the activities in the suffix. Our approach tackles this limitation by departing from the SM approach and, instead, training three separate models individually to predict the next activity, its inter-start time, and its processing time. This approach allows us to flexibly calculate waiting times and processing times separately.
@@
Other researchers have combined simulation with learning-based techniques. Camargo et al. [3] and Meneghello et al. [11] proposed integrating LSTM predictors with discrete-event simulation models derived from discovered process models. These methods simulate synthetic traces, including timestamps, under hypothetical “what-if” scenarios and can be used to evaluate process changes or capacity modifications. In contrast, our approach focuses on generating case suffixes for “as-is” processes by leveraging real-time event logs without assuming process changes or relying on synthetic generation.
@@ 3.1 Offline Phase
The offline phase (Fig.1) focuses on preparing the training data and learning predictive models necessary for case suffix prediction. This phase begins with transforming an event log into structured input sequences by extracting both intra-case features (capturing case-specific temporal patterns), and inter-case features (capturing system-wide dynamics, such as resource utilization and workload). These features are encoded and assembled into fixed-size sequences using an n-gram strategy to standardize inputs for model training. We train three specialized BiLSTM-based models – one each for predicting the next activity($\alpha$), inter-start time($\beta$), and processing time($\gamma$). Below, we describe the offline phase.
@@
3.1.1 Input:
To train models for case suffix prediction, we take as input a type of event log called an activity instance log. Table 1 contains an excerpt of such a log. Each row corresponds to an activity instance. For each activity instance, the proposed approach requires a case ID, activity name, start and end timestamps. Every row must have a value for each of these four attributes. However, for some activity instances (e.g., the second one in Case 2) the value of the end timestamp may be null (denoted as $\emptyset$). When an activity instance has a null end timestamp, it means that it has not yet completed. Our approach does not require other attributes in the log. In particular, it does not require a resource attribute nor a role attribute. On the other hand, the approach does assume that the case ID, activity name, start and end timestamps are accurate, otherwise the machine learning models would be trained to produce inaccurate case suffixes. Formally, we define an activity instance log as follows.
Definition 1(Activity Instance Log)
An Activity Instance Log $\mathcal{L}$, is a finite set of records $r\in\mathcal{L}$, where each record $r=(c,a,T^{\text{start}},T^{\text{end}})$ consists of the following elements: $c\in C$,where $C$ is the set of unique case identifiers for process instances; $a\in A$, where $A$ is the set of possible activities; and $T^{\text{start}},T^{\text{end}}\in\mathbb{T}\cup\{\emptyset\}$, representing the start timestamp and end timestamp of an activity, with $\mathbb{T}$ being the domain of time values and $\emptyset$ denotes a missing (null) timestamp.
@@ 3.1.1 Input:
Case ID Activity $T^{\text{start}}$ $T^{\text{end}}$
1 Received Query 08:00 08:05
Assigned to Rep 08:10 08:15
Query Resolved 08:20 09:00
Cust Notified 09:05 09:10
2 Received Query 09:15 09:20
Assigned to Rep 09:25 $\emptyset$
3 Received Query 10:00 10:05
Assigned to Rep 10:10 10:15
Query Resolved 10:20 11:00
@@ 3 Received Query 10:00 10:05
Definition 2(Activity Instance Trace of Log $\mathcal{L}$)
Given an activity instance log $\mathcal{L}$, an activity instance trace $\sigma_{c}$ of $\mathcal{L}$ is a sequence of activity instances $\sigma_{c}=\langle r_{c,1},r_{c,2},\dots,r_{c,n}\rangle$, such that:
  • All activity instances $r_{c,i}$ in $\sigma_{c}$ share the same case identifier $c$.
@@
Our approach takes as input an activity instance log $\mathcal{L}$ generated by a collection of cases over a specified time period and a time point called cutoff denoted by ($t_{\text{cutoff}}$). Given $\mathcal{L}{}$ and $t_{\text{cutoff}}$, we derive the training log, denoted as $\mathcal{L}_{train}$. The training log includes all $\sigma_{c}{}\in\mathcal{L}{}$ that fulfill the following two conditions:
@@
To train a deep learning model, we enhance each case in the training log with intra- and inter-case features. Alg. 1 describes the procedure for enhancing a given trace $\sigma_{c}{}$ with such features. This procedure is applied to each $\sigma_{c}{}$ $\in$ $\mathcal{L}_{train}$.
Input: An activity instance log $L$, An activity instance trace $\sigma_{c}\in L$
Output: An enhanced activity instance $\sigma_{c}$ including inter- and intra-case features
@@
2foreach trace prefix $P=\langle r_{c,1},\dots,r_{c,k}\rangle$ in $\sigma_{c}$ do
@@
4     Let $r_{c,i}=(c,a_{c,i},T^{\text{start}}_{c,i},T^{\text{end}}_{c,i})$;
@@ 10        $\delta^{\text{start}}_{c,i}\leftarrow T^{\text{start}}_{c,i}-T^{\text{start}}_{c,i-1}$;
11
12    $r_{c,i}\leftarrow r_{c,i}\oplus(\delta^{\text{proc}}_{c,i},\delta^{\text{start}}_{c,i})$
@@
14 Let $T^{\text{eval}}\leftarrow T^{\text{start}}_{c,k}$;
15
16 BusyResources $\leftarrow\emptyset$;
@@ 17  TotalActiveInstances $\leftarrow 0$;
18
19 foreach $r=(c^{\prime},a^{\prime},T^{\prime}_{\text{start}},T^{\prime}_{\text{end}})\in L$ do
20     if $T^{\prime}_{\text{start}}\leq T^{\text{eval}}\leq T^{\prime}_{\text{end}}$ then
21        Increment TotalActiveInstances;

### v2

# How Will My Business Process Unfold? Predicting Case Suffixes With Start and End Timestamps

@@ Abstract
Predictive process monitoring techniques support operational decision-making by predicting future states of ongoing cases in a business process. A subset of these techniques focuses on predicting the remaining sequence of activities for an ongoing case, known as case suffix prediction. Existing approaches for case suffix prediction typically generate sequences of activities with a single timestamp (e.g., the end timestamp). While this approach is useful in some contexts, it is insufficient for applications like resource capacity planning, where it is critical to reason about both the waiting time and processing time of each activity. Such information is necessary to accurately assess when resources will be engaged, optimize scheduling, and ensure capacity is adequate for future workloads. This paper introduces a technique for predicting case suffixes consisting of activities with both start and end timestamps. Specifically, the proposed technique predicts both the waiting time and processing time of each activity. Since the waiting time of an activity in a case depends on the availability of resources in other ongoing cases, the technique adopts a sweep-line approach, wherein the suffixes of all ongoing cases are predicted in lockstep, rather than predictions being made for each case in isolation. Additionally, the technique improves temporal accuracy by integrating resource availability information, automatically discovered from historical event logs, allowing for more realistic predictions in dynamic environments. An evaluation on real-life and synthetic datasets compares the accuracy of different instantiations of this approach, demonstrating the advantages of a multi-model approach to case suffix prediction over existing methods. The results highlight the improved performance in predicting both control-flow and temporal aspects of case suffixes, particularly in scenarios involving sparse and intermittent resource availability.
keywords:
Process Mining; Predictive Process Monitoring; Sequence Prediction
@@ 1 Introduction
Predictive Process Monitoring (PPM) plays a role in Business Process Management (BPM) by providing managers with data-driven projections on running cases, enabling more informed and effective decision making [28]. By leveraging models trained on event logs, PPM techniques offer runtime insights that predict the future states of ongoing process cases. For example, PPM techniques can predict the outcome of ongoing cases (e.g., will a customer accept or reject a product?) [34], the remaining time of ongoing cases [38], the next activity in a case, or the sequence of remaining activities in a case (a.k.a. the case suffix[28]. These predictions are useful for process optimization, resource planning, and decision-making, especially in operational environments where timely insights are needed to, for instance, maintain efficiency and to meet service level agreements.
Existing approaches for predicting the next activities in a process, based on its current state (a.k.a. case suffix), primarily focus on generating sequences of activities with a single timestamp (e.g., end timestamp) [6, 32, 12]. This is insufficient for use cases such as capacity planning and scheduling, where managers need both start and end timestamps to calculate processing and waiting times, assess resource utilization, and determine whether current capacity is adequate for the expected workload [3]. For instance, an operational manager may anticipate that, due to a flu epidemic, around 10% of the workforce in a specific business area will report sick next Monday morning. As a result, reduced capacity to handle tasks is expected throughout the day. What is needed is the projected workload on Monday morning, given the current set of cases and their current state. If predictions only provide the next activities between now and the end of next week, using only one timestamp, it becomes impossible to determine how many activity instances will be active during specific periods, such as between 8 AM and 10 AM. Without the start timestamp, it is not possible to discern whether activity instances are in progress, waiting, or recently completed. In contrast, if predictions provide both start and end timestamps for each activity, the number of activity instances active between 8 AM and 10 AM on Monday can be estimated. This enables informed decisions about the number of workers needed to manage that workload effectively.
Previous approaches primarily focus on the temporal and control flow aspects of process prediction, such as the sequence of next activities and activity timestamps. However, they typically overlook resource availability and utilization. Including these factors can directly enhance insights by allowing a better understanding of how activity durations are affected by resource constraints, how resources should be allocated, and how many activities can be executed concurrently. For example, if resources are limited during a specific period, the predicted duration of an activity may increase, or certain activities might be delayed, which in turn impacts the overall schedule. This gap in current approaches highlights an opportunity for improvement. By incorporating both the start and end timestamps of predicted activities, along with resource availability and utilization, managers can gain a more comprehensive view of the process. Such predictions would enable them to make more informed decisions about resource allocation, ensuring that tasks are completed efficiently even during periods of resource constraints.
@@
In this setting, we propose an approach for predicting case suffixes composed of activities with start and end timestamps. Our method uses a sweep-line-based technique [25], which incorporates resource availability features based on resource availability calendars, automatically discovered from historical event logs to predict case suffixes for all ongoing cases collectively, rather than individually. The approach follows a three-stage prediction process. First, a model predicts the next activity. Given this activity, a second model estimates the inter-start time, i.e., the time elapsed between the start of the previous activity and the start of the predicted activity. Finally, a third model predicts the processing time, i.e., the duration required to complete the activity once it has started. Using these predictions, we derive the start and end timestamps of each predicted activity instance in the case suffix. The start timestamp is obtained by adding the inter-start time to the start timestamp of the previous activity instance, while the end timestamp is derived by adding the processing time to the start timestamp of the predicted activity instance. This approach ensures accuracy by providing precise estimations of time intervals (inter-start time and processing time) that determine the timing of each activity. It is also complete in that it provides both start and end timestamps for each activity, offering a full timeline of the case suffix, which is essential for effective resource management, scheduling, and operational forecasting.
This article is an extended version of  [2]. The conference paper focuses on predicting case suffixes consisting of activities with start and end timestamps by employing a sweep-line-based approach. This article extends the conference paper by improving the temporal accuracy of the proposed approach by integrating information about resource availability based on resource availability calendars, automatically discovered from historical event logs. We hypothesize that this resource availability information, automatically discovered from historical event logs, will allow us to further refine our predictions and make them more robust and applicable to real-world scenarios, where resource constraints and availability influence process outcomes.
We conducted a two-pronged evaluation using synthetic and real-life event logs to assess the effectiveness of our multi-model approach. The synthetic evaluation examines how varying resource availability calendars affect the control-flow and temporal accuracy of case suffix prediction, simulating different on-off duty schedules, and resource constraints. This allows us to explore how the proposed approach adapts to changes in resource availability, which is a common challenge in real-world process management. Meanwhile, the evaluation with real-life logs compares our approach to existing multi-model approaches, using control-flow and temporal metrics to assess the accuracy and reliability of case suffix predictions in terms of both start and end timestamps.
The remainder of this article is structured as follows. Sec. 2 provides an overview of prior research related to case suffix prediction. Sec. 3 presents the proposed approach for case suffix prediction. Finally, Sec. 4 reports the empirical evaluation, and Sec. 5 concludes the paper and Sec. 6 outlines directions for future work.
@@ 2 Related Work
Predictive Process Monitoring (PPM) is a research stream within the broader field of Process Mining (PM) [9]. Unlike traditional descriptive approaches such as process discovery, conformance checking, and model enhancement, PPM incorporates predictive and monitoring capabilities that apply machine learning models to ongoing cases in a business process [37, 34]. By exploiting historical event data generated by organizational information systems, PPM transforms process analysis from a backward-looking to a forward-looking perspective [8]. Through this predictive capability, it becomes possible to estimate future activities, remaining execution time, or likely case outcomes, which are essential for data driven decision making [11, 8]. In practice, predictive models derived from PPM can support the timely allocation of resources [21], early identification of process bottlenecks [35], and adaptive interventions aimed at improving operational performance [7, 29].
In recent years, PPM research has increasingly focused on employing deep learning architectures to enhance the predictive performance of forecasting future activities within an ongoing case (a.k.a the case suffix prediction). Early studies such as those by Tax et al. [31] and Evermann et al. [10] introduced LSTM-based models to predict the next activity, case suffix, and remaining time until completion, using one-hot and embedded sequence encodings. Camargo et al. [6] extended these efforts through multi-output LSTM models that jointly predict activity, end timestamp, and role. Other approaches have leveraged architectures, such as GANs [33], transformers and attention-based encoder-decoder models [39, 27], and reinforcement learning-based samplers like DOGE [26]. Pasquadibisceglie et.al., [23] also explored CNNs and, more recently, large language models (LLMs) [24] for case suffix prediction.
@@
The above approaches follow a single model (SM) paradigm, which jointly predicts the next activity and one timestamp (either the start or the end time). None of these approaches is designed to estimate both the start and the end timestamps of the activities in the case suffix. In other words, none of these approaches separates the waiting time and the processing time of the activities in the suffix. Our approach tackles this limitation by departing from the SM approach and, instead, training three separate models individually to predict the next activity, its inter-start time, and its processing time. This approach allows us to flexibly calculate waiting times and processing times separately.
@@
Furthermore, researchers have explored the integration of learning-based predictors with simulation models. Camargo et al. [4] and Meneghello et al. [22] combined LSTM predictors with discrete-event simulation derived from discovered process models to generate synthetic traces and evaluate “what-if” scenarios. These hybrid methods are valuable for studying process changes or capacity adjustments. In contrast, our approach focuses on generating case suffixes for “as-is” processes by leveraging real-time event logs without assuming process changes or relying on synthetic generation.
Recently, a few of the studies have focused on integrating resource-related features to enhance both predictive accuracy and model interpretability [16, 18, 15, 36]. Klijn et al. [16] investigated resource waiting-time patterns to better understand process performance, showing that analyzing behavior at the resource level yields more meaningful and interpretable insights than relying on aggregated waiting-time averages. Similarly, Kunkler et al. [18] demonstrated that considering the time individual resources require to complete tasks can outperform conventional allocation strategies, underscoring the role of human and resource characteristics in process optimization. Additionally, a recent survey by Ali et al. [3] explored the various causes and interpretations of waiting times related to resource contention and unavailability, aiming to better understand process bottlenecks through the analysis of resource behavior.
Beyond predictive modeling, several studies have revisited business process simulation (BPS) from a resource-oriented standpoint. For example, the authors in [15] proposed a resource-first simulation approach that constructs multi-agent systems from event logs, effectively capturing heterogeneous resource behaviors and interactions. Their findings indicated improved accuracy, reduced computational overhead, and greater interpretability across various process settings. In the same direction, Tour et al. [36] introduced Agent Miner, an algorithm designed to discover agent interaction models directly from event data, resulting in representations of business processes that more closely mirror real-world execution compared to traditional discovery techniques.
Despite recent advancements in case suffix prediction using PPM, most approaches remain focused primarily on control-flow aspects [17]. A recent benchmark and review by Rama et al. [28] highlights that newer techniques, such as LSTMs and other deep learning models, continue to adopt a case-specific control-flow perspective, often overlooking the influence of resource availability factors. While previous studies have incorporated resource-related information, such as the role or resource performing an activity, no study to date has systematically integrated resource availability aware features such as working shifts, scheduled breaks, workload conditions, or expected availability times into predictive models for the case suffix prediction problem. Collectively, these gaps suggest that integrating resource availability aware information could enhance both the accuracy and performance of process mining tasks. Building on this foundation, our study advances the PPM literature by approaching the case suffix prediction problem from a resource-centric perspective, bridging predictive performance with the analysis of resource availability aware behavior for more accurate predictions of case suffixes in an ongoing business process.
@@ 3.1 Offline Phase
The offline phase (Fig.1) focuses on preparing the training data and learning predictive models necessary for case suffix prediction. This phase begins with transforming an event log into structured input sequences by extracting intra-case features (capturing case-specific temporal patterns), resource contention features (capturing system-wide dynamics, such as resource utilization and workload) and resource availability features. These features are encoded and assembled into fixed-size sequences using an n-gram strategy to standardize inputs for model training. We train three specialized BiLSTM-based models – one each for predicting the next activity($\alpha$), inter-start time($\beta$), and processing time($\gamma$). Below, we describe the offline phase.
@@
3.1.1 Input
To train models for case suffix prediction, we take as input a type of event log called an activity instance log. Table 1 contains an excerpt of such a log. Each row corresponds to an activity instance. For each activity instance, the proposed approach requires a case ID, activity name, a resource, start and end timestamps. Every row must have a value for each of these five attributes. However, for some activity instances (e.g., the second one in Case 2) the value of the end timestamp may be null (denoted as $\emptyset$). When an activity instance has a null end timestamp, it means that it has not yet completed. On the other hand, the approach does assume that the case ID, activity name, start and end timestamps are accurate, otherwise the machine learning models would be trained to produce inaccurate case suffixes. Formally, we define an activity instance log as follows.
Definition 3.1 (Activity Instance Log).
An Activity Instance Log $\mathcal{L}$, is a finite set of records $e\in\mathcal{L}$, where each record $e=(c,a,r,T^{\text{start}},T^{\text{end}})$ consists of the following elements: $c\in C$, where $C$ is the set of unique case identifiers for process instances; $a\in A$, where AA is the set of possible activities; $r\in R$, where $R$ is the set of resources executing the activity; and $T^{\text{start}},T^{\text{end}}\in\mathbb{T}\cup\{\emptyset\}$, representing the start timestamp and end timestamp of an activity, with $\mathbb{T}$ being the domain of time values and $\emptyset$ denotes a missing (null) timestamp.
@@ 3.1.1 Input
Case ID Activity Resource $T^{\text{start}}$ $T^{\text{end}}$
1 Received Query System 08:$0$ 08:05
Assigned to Rep Supervisor 08:10 08:15
Query Resolved Rep A 08:20 09:$0$
Cust Notified System 09:05 09:10
2 Received Query System 09:15 09:20
Assigned to Rep Supervisor 09:25 $\emptyset$
3 Received Query System 10:$0$ 10:05
Assigned to Rep Supervisor 10:10 10:15
Query Resolved Rep B 10:20 11:$0$
@@ 3 Received Query System 10:$0$ 10:05
Definition 3.2 (Activity Instance Trace of Log $\mathcal{L}$).
Given an activity instance log $\mathcal{L}$, an activity instance trace $\sigma_{c}$ of $\mathcal{L}$ is a sequence of activity instances $\sigma_{c}=\langle e_{c,1},e_{c,2},\dots,e_{c,n}\rangle$, such that:
  • All activity instances $e_{c,i}$ in $\sigma_{c}$ share the same case identifier $c$.
@@
Our approach takes as input an activity instance log $\mathcal{L}$ generated by a collection of cases over a specified time period and a time point called cutoff denoted by ($t_{\text{cutoff}}$). Given $\mathcal{L}{}$ and $t_{\text{cutoff}}$, we derive the training log, denoted as $\mathcal{L}_{train}$. The training log includes all $\sigma_{c}{}\in\mathcal{L}{}$ that fulfil the following two conditions:
@@
To train a deep learning model, we enhance each case in the training log with resource contention, intra-case and resource availability features. Alg. 1 describes the procedure for enhancing a given trace $\sigma_{c}{}$ with such features. This procedure is applied to each $\sigma_{c}{}$ $\in$ $\mathcal{L}_{train}$.
Input: An activity instance log $L$, an activity instance trace $\sigma_{c}\in L$, resource calendars $C_{r}$ for all resources
Output: An enhanced activity instance trace $\sigma_{c}$ including intra-, $R_{\text{cont}}$, and $R_{\text{avail}}$ features
@@
2foreach trace prefix $P=\langle e_{c,1},\dots,e_{c,k}\rangle$ in $\sigma_{c}$ do
3
@@
5      Let $e_{c,i}=(c,a_{c,i},T^{\text{start}}_{c,i},T^{\text{end}}_{c,i})$;
@@ 12       $\delta^{\text{start}}_{c,i}\leftarrow T^{\text{start}}_{c,i}-T^{\text{start}}_{c,i-1}$;
14      end if
15    $e_{c,i}\leftarrow e_{c,i}\oplus(\delta^{\text{proc}}_{c,i},\delta^{\text{start}}_{c,i})$;
16
17   end for
18
19  Let $T^{\text{current}}\leftarrow T^{\text{start}}_{c,k}$;
20
21 BusyResources $\leftarrow 0$;
@@ 22 TotalActiveInstances $\leftarrow 0$;
23
24 foreach $e=(c^{\prime},a^{\prime},T^{\prime}_{\text{start}},T^{\prime}_{\text{end}})\in L$ do
25    if $T^{\prime}_{\text{start}}\leq T^{\text{current}}\leq T^{\prime}_{\text{end}}$ then
26       TotalActiveInstances $\leftarrow$ TotalActiveInstances +1+1;